Submit your detailed Question here...if ax^2 + bx +6 = 0 does not have 2 distinct real roots, , then least value of 3a + b is?

Grade:12

3 Answers

vinay kant
11 years ago

Dear Sijin,

According to given condition,

D<0;

b^2 - 24a <0

b^2 < 24a

as it does not have real roots thus

its discriminant is equal to zero

i.e.         b^2 - 4*a*6= b^2 - 24a=0

          a = b^2 / 24

     now take  y=3a+b

put value of a and differentiate it and you will get

dy/da=b/4 +1=0

b= -4

is the minimum value.............


    

Sajin SS
11 years ago

Yes..after that?

mycroft holmes
11 years ago